OpenStudy (anonymous):

Medal A pharmaceutical company combined 485,100 mg of acetaminophen with 9,900 mg of oxycodone to fill an order from a pharmacy for a powerful prescription pain-killer. The pain-killers come in 550 mg tablets. How many tablets were created by the company for this shipment? ? tablets

OpenStudy (snowsurf):

Eat tablet will have 550 mg and total = 485100 + 9900 = 495,000 So 495000/550 = 900. So they can produce 900 tablets.